OpenStudy (anonymous):

Let a>0 (a). Show using L'Hospitals rule that: \[\lim_{x \rightarrow 0} x ^{a}\ln(x)=0\] (b). By setting x=ln(t) in (a). show that: \[\lim_{x \rightarrow -\infty} |x|^{a} e ^{x}=0\] I need help with (b). I don't get the question, how do I show that (b). is true by setting x=ln(t) in (a). it doesn't make any sense to me.

OpenStudy (sirm3d):

the proposed substitution

OpenStudy (sirm3d):

\[x=\ln t\] means that \[t=e^x\]

OpenStudy (anonymous):

I've got that but how can I use it?

OpenStudy (sirm3d):

as \[x \rightarrow -\infty, t \rightarrow ?\]

OpenStudy (anonymous):

Zero?

OpenStudy (anonymous):

e^-infinty = 0

OpenStudy (sirm3d):

correct

OpenStudy (sirm3d):

so we shift from

OpenStudy (sirm3d):

\[\lim_{x \rightarrow -\infty} \] to \[\lim_{t \rightarrow 0^+}\]

OpenStudy (sirm3d):

in terms of t, \[\left| x \right|^a = ?\]

OpenStudy (anonymous):

\[|\ln (t)|^{a}\] ?

OpenStudy (sirm3d):

that's right. and what about \[e^x\]

OpenStudy (anonymous):

We could set it as just t, right?

OpenStudy (sirm3d):

right

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0} |\ln(t)|^{a}t\]

OpenStudy (anonymous):

Because t->0 shouldn't the whole expression be equal to 0 then?

OpenStudy (sirm3d):

not necessarily

OpenStudy (sirm3d):

what about \[\left| \ln t \right|^a\]

OpenStudy (anonymous):

\[\ln(t) \rightarrow -\infty\] but since it's the abolute value its just \[\infty \]

OpenStudy (sirm3d):

\[\left| \ln t \right|^a \rightarrow +\infty \]. you are correct

OpenStudy (sirm3d):

the transformed problem is an indeterminate form \[0*\infty \] which should be resolved either as \[\frac{ 0 }{ 0 }\] or \[\frac{ \infty }{ \infty }\]

OpenStudy (anonymous):

Right, so I could tranform it to the right form and use L'Hospitals rule to get the answer correct

OpenStudy (anonymous):

Should I use the form \[\frac{ t }{ |\ln(t)|^{-a} }\] and then go for L'Hospital?

OpenStudy (sirm3d):

how about the other form?

OpenStudy (anonymous):

Just as it was, is it possible to use L'hostial on multiplication form?

OpenStudy (sirm3d):

lhopitals rule must be in quotient form

OpenStudy (anonymous):

Which form are you thinking about then?

OpenStudy (sirm3d):

i'm trying your quotient form

OpenStudy (anonymous):

\[-\frac{ a|\ln(t)|^{-a-1} }{ t }\] Can that be the correct derivation? If so it should give us the right answer \[\lim_{t \rightarrow 0}\frac{ 1 }{ -\frac{ a|\ln(t)|^{-a-1} }{ t } } = \frac{ 1 }{- \infty }=0\] Can this be right?

OpenStudy (sirm3d):

we'll look at the 2nd quotient form \[\frac{ \left| ?\ln t \right|^a }{ \frac{ 1 }{ t } }\], noting that a>0

OpenStudy (sirm3d):

since t->0+, \[\left| \ln t \right|=-\ln t \]

OpenStudy (anonymous):

Right

OpenStudy (sirm3d):

the indeterminate form is inf/inf. so we'll apply l'hop's rule once

OpenStudy (sirm3d):

i hate getting disconnected in the middle of the equation

OpenStudy (anonymous):

\[\frac{ \frac{ -a \ln ^{a-1}(t) }{t } }{\frac{ -1 }{ t ^{2} } }\]

OpenStudy (sirm3d):

right. when simplified becomes \[\frac{ a(-\ln t)^{a-1} }{ 1/t }\]

OpenStudy (sirm3d):

here's the catch to the problem

OpenStudy (sirm3d):

\[(-\ln t)^a \rightarrow +\infty \] for as long as the exponent a > 0.

OpenStudy (anonymous):

So it's still \[\frac{ \infty }{ \infty } \]

OpenStudy (sirm3d):

right. but sooner or later, the exponent becomes negative after successive applications of l'hopital's rule

OpenStudy (anonymous):

\[\frac{ (a-1)a(-\ln(t))^{a-2} }{ 1/t }\]

OpenStudy (sirm3d):

after k applications of the rule, the indeterminate form is now \[a(a-1)(a-2)...(a-k+1)\frac{ (-\ln t)^{a-k} }{ 1/t }\] where a-k is negative

OpenStudy (sirm3d):

so that \[(-\ln t)^{a-k} \rightarrow 0\] because a-k is negative

OpenStudy (sirm3d):

the quotient form is now \[a(a-1)(a-2)...(a-k+1)\frac{ 0 }{ \infty }\] and the problem is solved!

OpenStudy (anonymous):

Oh, that's clever, I get it

OpenStudy (anonymous):

Thank you so much for you're help I can't thank you enought this one has been on my mind for a couple of days now!